OpenStudy (anonymous):

solve and show work 4(c-3)-6=10 c=?

OpenStudy (anonymous):

multiply 4(c-3) at first

OpenStudy (inkyvoyd):

The, isolate the variable.

sam (.sam.):

\[\begin{array}{l} 4 (c-3)-6=10 \\ \text{Add }6\text{ \to both sides:} \\ 4 (c-3)=16 \\ \text{Divide both sides by }4: \\ c-3=4 \\ \text{Add }3\text{ \to both sides:} \\ c=7 \\\end{array}\]

OpenStudy (inkyvoyd):

@.Sam. , have fun texing.

OpenStudy (inkyvoyd):

Wow. posted before I finished.

sam (.sam.):

lol